OpenStudy (anonymous):

a metal strip is being installed around a workbench that is 12 ft long and 2 ft wide. if the stripping costs $7 per foot, find the total cost of the stripping.

OpenStudy (anonymous):

if it cost 7 dollars for one foot how much will it cost for 2 feet?

OpenStudy (anonymous):

this problem is worded weird. usually you buy things is ft^2

OpenStudy (anonymous):

168 $

OpenStudy (anonymous):

I think its $98

OpenStudy (anonymous):

you need to work out the perimeter of the workbench first. P= 2(l+w) the length is 12ft and the width is 2 ft, so P= 2(12ft + 2ft) = 28ft so now we know the perimeter (28ft) and it's $7 per feet. 28ft x $7 = $196

OpenStudy (anonymous):

woops i found half of the perimeter sorry.

OpenStudy (anonymous):

i thought the area of rectangle.

OpenStudy (anonymous):

I assume: $7/ ft^2, then: $7 * ( 12 * 2) = $ 168